Dorcan Corporation manufactures and sells T-shirts imprinted with college names and slogans. Last year, the shirts sold for $9.00 each, and the variable cost to manufacture them was $2.25 per unit. The company needed to sell 21,000 shirts to break-even. The after tax net income last year was $5,340. Donnelly's expectations for the coming year include the following: (CMA adapted)

  • The sales price of the T-shirts will be $14.
  • Variable cost to manufacture will increase by one-third.
  • Fixed costs will increase by 10%.
  • The income tax rate of 40% will be unchanged.


The selling price that would maintain the same contribution margin ratio as last year is:

  • $11.50.

  • $10.25.

  • $12.00.

  • $11.00.
Answer #1

Selling price per unit = $9

Variable cost per unit = $2.25

Contribution margin per unit = Selling price per unit-Variable cost per unit

= 9-2.25

= $6.75

Contribution margin ration = Contribution margin / Selling price per unit

= 6.75/9

= 75%

For coming year

Variable cost per unit will increase by 1/3

Increase in variable cost per unit = 6.75 x 1/3

= $2.25

Variable cost per unit = 6.75+2.25

= $9

Contribution margin ratio = Contribution margin/selling price per unit

0.75 = ( Selling price per unit - variable cost per unit) / selling price per unit

0.75 = ( Selling price per unit - 9)/Selling price per unit

0.75 Selling price per unit = Selling price per unit -9

Selling price per unit = 9/0.75

= $12

The selling price that would maintain the same contribution margin ratio as last year is $12.

Third option is correct option.

Kindly comment if you need further assistance.

Thanks‼!